3. Find the particular solution of the linear equation dy/dx + 1/x y = x if y(1) = 0


1
Expert's answer
2022-02-02T15:15:33-0500
"y'+\dfrac{1}{x}y=x"

Integration factor


"\mu(x)=e^{\int(dx/x)}=x"

"xy'+y=x^2"

"d(xy)=x^2 dx"

Integrate


"\int d(xy)=\int x^2 dx"

"xy=\dfrac{x^3}{3}+C"

"y=\dfrac{x^2}{3}+\dfrac{C}{x}"

Given "y(1)=0"


"0=\dfrac{(1)^2}{3}+\dfrac{C}{1}=>C=-\dfrac{1}{3}"

The solution of the initial value problem is


"y=\dfrac{x^2}{3}-\dfrac{1}{3x}"